# Simplifying Triangle Meshes in the Wild

### v1

@@ 9.3. Perceptual User Studies
Besides all qualitative and quantitative evaluations, we perform a user study to gather perceptual opinions on mesh simplification results. Specifically, the participants were told to perform the user study in the context of a common online game scenario where the players are using low-end devices (e.g., mobile phones) to play online video games with other players. Due to the high performance requirement (e.g., frame rates, bandwidth), the 3D assets in the game have to satisfy a fixed budget (e.g., the total number of vertices in the scene). Under this situation, we asked participants to fill out an online survey (see inset) about their preferences in a variety of simplified 3D assets obtained from our method, the probabilistic quadric error simplification [Trettner and Kobbelt, 2020], and our invented baseline [G97+L98+T20] mentioned in section 4. The results from these methods are presented in an arbitrary order in the survey to avoid order bias. We invited 69 participants (51 male, 17 female, 1 non-binary individual, age ranging from 11-50, 75% graduate students and 25% working professionals including engineers and artists) to conduct the user study. Each of them was asked to select their favorite mesh simplification results out of the 10 user-created meshes presented in the paper (see Figure 22 for the collection) representing a wide range of meshes from organic shapes, 3D scenes, and man-made objects. From the total 690 responses (69$\times$10), 82.5% of them favored our results, 10.4% favored the results from [Trettner and Kobbelt, 2020], and 7.1% favored [G97+L98+T20]. The result demonstrates that our method leads to consistently better results across different meshes and individuals.
10. Discussion & Conclusion
We present a practical mesh simplification method that can achieve better results on human-created meshes that commonly appear in practice. We have also described an effective solution for copying the surface attributes without impacting the quality of the mesh simplification process. Our method pushes the robustness of mesh simplification to gracefully handle any triangle meshes, including non-manifold meshes and meshes with multiple connected components. This enables smoother performance by reducing the complexity of 3D models without compromising visual quality, enabling fast rendering and interactive simulation for real-time applications like online gaming, augmented reality, and virtual reality, especially on low-power devices such as mobile phones and VR headsets, where computational resources are limited.
Despite being robust, our method does not guarantee that the copied texture data comes from the out-most parts of the surface. Therefore, when simplifying meshes with multiple layers, the resulting texture can contain data from interior layers, which can be undesirable (see Figure 21).
Future work can address this problem by considering external visibility information while computing the mapping from the final mesh to the input surface. Our research contributes to the field of robust geometry processing and underscores the need to further enhance the robustness of downstream geometric algorithms, particularly because the output of our method, like the input, may contain defects such as non-manifold structures.

### v2

@@ Appendix D User Study Details
We performed two user studies to gather perceptual opinions on our mesh simplification results. Specifically, the participants were instructed to perform the user study within the context of a common online game scenario where the players are using low-end devices (e.g., mobile phones) to play online video games with other players. Due to the high performance requirements (e.g., frame rates, bandwidth), the 3D assets in the game have to satisfy a fixed budget (e.g., the total number of vertices in the scene). Under this situation, we asked participants to fill out online surveys about their preferences among a variety of simplified 3D assets obtained from our method and baselines.
One of the studies evaluated the quality of textured mesh simplification. For each response, we randomly selected a mesh from the PolyHaven dataset, simplified with our method and the textured QEM by Garland and Heckbert [1998], and asked the participant to select their preferences among (1) our method, (2) [Garland and Heckbert, 1998], and (3) comparable (see Figure 24 for some examples in the dataset). We included comparable as one of the options because our method “converges” to QEM when the input is a closed manifold mesh and this can avoid users randomly picking one out of comparable outputs. The results from the two methods are presented in an arbitrary order in the survey to avoid the order bias. From the 591 responses we collected, 81.9 % of them indicated that our method is either comparable or outperforming the baseline, suggesting that our method leads to better simplification results (see section 5 for more details).
[…]
Figure 25. We conducted a user study to evaluate the perceptual preference across different mesh simplification techniques, including (1) ours, (2) [Trettner and Kobbelt, 2020], and (3) [G97+L98+T20] a combination of [Trettner and Kobbelt, 2020; Garland and Heckbert, 1997; Lindstrom and Turk, 1998]. The figure contains all the meshes used in the user studies. Note that in the user study, the order of different methods is randomly permuted to avoid order bias, but in this figure we display them in a consistent order for clarify.
For these two studies, we invited 73 participants (51 male, 21 female, 1 non-binary individual, age ranging from 11-50, 75% graduate students and 25% working professionals including engineers and artists) to conduct the user study.